OpenStudy (anonymous):

2(x +7)^2 = 64 what is x?

OpenStudy (anonymous):

You need to isolate x. So start by dividing by 2.

OpenStudy (anonymous):

that would be (x + 7) = 32 rite?

OpenStudy (anonymous):

It would be \[(x+7)^2=32\]

OpenStudy (anonymous):

yeah ok

OpenStudy (anonymous):

so where do i go from there>

OpenStudy (anonymous):

Well you want to get x by itself and it's in a quantity that is squared. So you need to undo the square...

OpenStudy (anonymous):

would that be \[x + 7 = 4\sqrt{2}\]

OpenStudy (anonymous):

You bet :) Now just get x by itself!